# Speech emotion recognition from voice messages recorded in the wild

### v1

@@ IV-C Pre-trained model embeddings and eGeMAPS
V Results
The cross-validation results, depicted in Fig. 2, provide a comprehensive overview of the performance across the different databases and emotion labels. In this visualization, the unweighted accuracy of the optimal baseline model using acoustic features and machine learning (called eGeMAPS) is indicated in gray. Adjacent to this baseline, the results for all models employing pre-trained models (named Embeddings) are displayed, along with a third column with the method integrating the speech embeddings extracted from pre-trained models and the eGeMAPS features together (called Emb+eGeMAPS). The figure indicates the mean and standard deviation of the UA across the five folds (four folds in the case of IEMOCAP).
Figure 2: Cross-validation results for the three methods implemented (eGeMAPS, Embeddings and Emb+eGeMAPS) across the different databases (EMOVOME, IEMOCAP and RAVDESS) and emotion labels (valence, arousal and categories of emotions).
Regarding the baseline eGeMAPS models, the lowest UA values are from the EMOVOME database, with 41-44% for valence, 41-51% for arousal and 31.42% for four categories. IEMOCAP achieves 50.20%, 53.21% and 59.50% for valence, arousal and categories, respectively. RAVDESS achieves the highest UA results for valence and arousal, with 57.19% and 62.48%, respectively. For the eight categories, it obtains 45.52%. Overall, SVC was preferred to KNN in most cases.
@@ V Results
Finally, the Emb+eGeMAPS method obtains similar results to the previous approach. EMOVOME achieves similar results in both valence and arousal (53-59% and 52-60% respectively), while the UA in 4-emotions classification decreases (40.22%). A possible reason is that EMOVOME was recorded in-the-wild conditions. Therefore, eGeMAPS features may be affected by microphone quality and background noise [52]. Conversely, IEMOCAP and RAVDESS were recorded in a controlled environment, and the UA values increased around 1-3% compared to the previous approach (except for the emotion categories with IEMOCAP, which slightly decreased). IEMOCAP achieves 62.48%, 60.41% and 68.04% for valence, arousal and categories, respectively. RAVDESS obtains 72.70%, 75.27% and 69.58% for valence, arousal and categories, respectively. Again, unispeech-L is the best option in the majority of cases, but now the other pre-trained models, which exhibited lower results in the previous approach, demonstrated significant improvement when combined with the eGeMAPS features, particularly in the cases of w2v2-xlsr-53 and w2v2-L-robust.
Using the hyperparameters of the models obtaining the highest UA in cross-validation, we trained new models using the development set and evaluated them on the test set. The results are shown in Table V and VI for valence and arousal prediction, respectively. The tables include the model achieving the highest cross-validation results among the three implemented methods: eGeMAPS, Embeddings and Emb+eGeMAPS. The test results follow the trends found in cross-validation, with RAVDESS outperforming IEMOCAP and EMOVOME across the two dimensions (73.54% in valence and 71.94% in arousal). IEMOCAP UA scores are higher than EMOVOME scores in arousal prediction (61.20% vs. 43.57-58.73% respectively), whereas, for valence prediction, IEMOCAP and EMOVOME achieve similar results (60.40% for the former and 57.53-61.64% for the latter). In the case of EMOVOME, we also explored the difference between expert (E) and non-expert (N) annotations, as well as their combination (C). Table V and VI show a comparison of results among them in the EMOVOME database, and their confusion matrix is represented in Fig. 3.
TABLE V: Test results for 3-class valence prediction. “PTM” specifies the pre-trained model applied. For EMOVOME, the type of rater is indicated: expert (E), non-expert (N), or combined (C).
Database Method PTM WA UA
@@
TABLE VI: Test results for 3-class arousal prediction. “PTM” specifies the pre-trained model applied. For EMOVOME, the type of rater is indicated: expert (E), non-expert (N) or combined (C).
Database Method PTM WA UA
@@
Figure 3: Confusion matrix for the test samples for valence and arousal prediction on EMOVOME. The rater is indicated: expert (E), non-expert (N) or combined (C).
An important limitation in IEMOCAP and RAVDESS is the label transformation applied to obtain the valence and arousal categories for their comparison with EMOVOME. For IEMOCAP, we selected two thresholds to categorize the continuous valence and arousal scores into the three categories studied. The model misclassifies those samples close to the thresholds, as shown in Fig. 4, where the bars indicate the distribution of the original valence and arousal values in IEMOCAP for the test samples, and the legend refers to the categories formed by converting labels. In the case of RAVDESS, we used Russell’s circumplex model of affect to derive the valence and arousal categories. Fig. 5 shows the distribution of the original emotion categories in RAVDESS for the test samples, the darker colors indicate the misclassified samples, and the legend refers to the categories resulting from the transformation of labels.
Figure 4: Prediction errors in the test for valence and arousal in IEMOCAP. The darker colors indicate the misclassified samples.
Figure 5: Prediction errors in test for valence and arousal in RAVDESS. The darker colors indicate the misclassified samples.
Considering the emotion categories, test results for the three databases are presented in Table VII. It includes the model achieving the highest CV results among the three implemented methods. RAVDESS obtains the highest UA score (75.00%), followed by IEMOCAP (69.58%) and EMOVOME (42.58%). To evaluate which emotions are misclassified, Fig. 6 shows the confusion matrix for each database.
TABLE VII: Test results for emotion categories prediction.
@@
Figure 7: Evaluation of gender fairness for the three labels and databases.
VI Discussion
In this work, we have developed speech emotion recognition models for EMOVOME, a natural database comprising spontaneous emotions from real voice messages collected in the wild. We have compared different methodologies to create the SER models, and we have explored the influence of annotators’ labels and the speaker’s gender on their performance. Additionally, we have compared the results with two other reference databases in the literature, IEMOCAP and RAVDESS. The UA is used for comparison in all cases since it is more suitable for assessing unbalanced data. The subsequent subsections individually address each of the research questions, and we finally discuss the limitations and future research directions.
VI-A RQ1: What is the performance of SER models in EMOVOME when utilizing classical acoustic features versus state-of-the-art speech embeddings?
@@ VI Discussion
We compared three methods used to create the SER models: the baseline using acoustic features and machine learning (eGeMAPS), the pre-trained models (Embeddings) and the integration of pre-trained models and eGeMAPS features (Emb+eGeMAPS). As shown in Fig. 2, overall, the embedding approach significantly improves model performance compared to the non-transformer baseline, leading to an approximately 10% improvement in unweighted accuracy across all combinations. Unispeech-L consistently showcased better results than other pre-trained models (as shown in [45]), closely followed by hubert-L (which also outperforms wav2vec2 models in previous studies [11]). Interestingly, for the EMOVOME database, these two pre-trained models are followed by w2v2-xlsr-53-spa for valence and categories prediction and by x-vectors for arousal prediction. For the English databases, w2v2-xlsr-53-spa was surpassed by w2v2-xlsr-128 and x-vectors, and in general, the models pre-trained on multiple languages performed worse than those trained on English-only data (as in [11]). The lowest scores for all databases correspond to w2v2-xlsr-53 and w2v2-L-robust. The former also obtained the worst results in [63] for a multilingual model with Spanish and English among several pre-trained multilingual models. As for w2v2-L-robust, it obtained the best results on IEMOCAP in [11], but the model was trained on the MSP-Podcast corpus and was only tested on IEMOCAP. As for the Emb+eGeMAPS method, it shows similar CV results to the embeddings for the EMOVOME database, and UA values slightly increase (max. 3%) for IEMOCAP and RAVDESS in some cases. One potential explanation is that EMOVOME was recorded in natural, uncontrolled conditions, which might impact the reliability of eGeMAPS features, unlike the other two databases that were conducted in a controlled environment. Unispeech-L is again the top choice in most cases, but notably, other pre-trained models that initially performed lower (e.g. w2v2-xlsr-53 and w2v2-L-robust) show significant improvement when combined with eGeMAPS features. Test results mirror cross-validation trends, with RAVDESS consistently outperforming IEMOCAP and EMOVOME across the three predictions. In arousal prediction, IEMOCAP UA scores are higher than EMOVOME, while similar results are observed for valence prediction between IEMOCAP and EMOVOME.
Considering the test results (see Tables V, VI, VII), they follow the trends found in cross-validation, obtaining similar performance scores. In summary, we found that Embeddings and Emb+eGeMAPS give similar results for the top-performing pre-trained model, i.e., Unispeech-L. Both approaches improve the speech baseline models presented in [29] in approximately 10% UA for the combined label in valence (61.64% vs 49.27%) and arousal (55.57% vs 44.71%). These results also improve the test UA score obtained in [29] using the transcriptions for arousal (47.43%). Nevertheless, our speech-based models for valence do not surpass the transcription-based model presented in [29] (61.15%). This suggests that both speech and text modalities offer complementary information across various emotion dimensions, underscoring the benefit of adopting a dimensional model approach for emotions.
VI-B RQ2: To what extent does the performance of EMOVOME, a natural database, compare with elicited and acted reference databases in the literature, such as IEMOCAP and RAVDESS?
We compared the results of our natural database, EMOVOME, with other two databases of different natures. We selected the well-known IEMOCAP database, including elicited speech, and RAVDESS, comprising acted recordings. We compared them in valence and arousal dimensions (see Tables V and VI), and emotion categories (see Table VII).
@@
Considering the prediction of emotion categories, the test results follow the trend in cross-validation, where EMOVOME obtains the lowest evaluation metric (45.58% UA), followed by IEMOCAP (69.58% UA) and finally RAVDESS (75.00% UA) (even though the first two classify 4 emotions and the later 8 emotions). Although IEMOCAP contains elicited speech, part of the data consists of actors performing scripted dialogues, which could lead us to expect higher classification results. However, there is a notable difference between IEMOCAP and RAVDESS. The former contains utterances whose text content is different throughout the database, i.e., it is text-independent (same as EMOVOME). Conversely, RAVDESS is text-dependent, as the actors portrayed emotions using two fixed sentences. This could have caused the models for RAVDESS to prioritize variations related to emotion rather than differences in semantic content. Consequently, there is a performance gap in IEMOCAP compared to the other acted databases, as highlighted in [64]. Nevertheless, our classification results (70.59% WA, 69.58% UA) are comparable to the state of the art for IEMOCAP, which is in the range from 60.0% to 74.3% UA [11], especially considering that we implemented a speaker-independent approach, unlike other previous studies. Additionally, we examined the misclassified emotions for each database in Fig. 6. For EMOVOME, surprise is not correctly predicted in any case and is mainly confused with happy and neutral emotion, but this category is underrepresented in the data (see Table III). In the case of IEMOCAP, all emotions are mainly mistaken for the neutral category. For RAVDESS, the majority of emotions are accurately classified ($>$72%), except for happy, which is sometimes misclassified as angry, and sad, which is mistaken for calm.
@@
Regarding the test results for arousal and valence dimensions, the SER models for the EMOVOME database achieve UA values of 61.64% for valence and 55.57% for arousal, considering the combined label between expert and non-experts. Surprisingly, the IEMOCAP database obtains similar values to EMOVOME, particularly 60.04% for valence and 61.20% for arousal. As for RAVDESS, the UA values are 73.54% and 71.94% for valence and arousal, respectively. Initially, one might have anticipated better outcomes in valence, given that earlier studies [65, 11] found that pre-trained models inherently capture linguistic information in the audio signal, aiding valence prediction. However, in our approach, we utilized pre-trained embeddings solely as feature extractors to obtain speech embeddings without fine-tuning the transformer layers for SER. This step has proven to be fundamental for the models to effectively learn the semantic content [11]. Furthermore, the data used for pre-training significantly impacts the models’ capacity to capture linguistic information, with the inclusion of multi-lingual data adding complexity to the task [11]. These considerations might explain the relatively minor differences observed between valence and arousal predictions. In the case of RAVDESS, the use of fixed semantic content during recording prevented pre-trained models from leveraging text information for valence prediction. Furthermore, an important limitation in both IEMOCAP and RAVDESS is the label transformation applied to obtain the valence and arousal categories, as well as the unbalanced distribution of samples (see Table III). In the case of IEMOCAP, errors in the test occur mostly around the thresholds selected to obtain the categories (see Fig. 4). Moreover, the arousal data is highly imbalanced, primarily dominated by samples in the neutral category. However, for valence, there is a relatively more balanced distribution across the three categories. For RAVDESS, in valence prediction, the majority of errors are related to the happy and calm category (see Fig. 5). Calm was categorized as positive, but it is close to the neutral valence. This ambiguity may have influenced the model’s positive valence predictions, especially considering these samples constitute half of the training data of this class. In arousal prediction, the model tends to make numerous errors in predicting disgust and sad. While disgust was labeled as neutral arousal, its proximity to neutral activation allows for variations, contributing to prediction inaccuracies. The same applies to sadness, as it is also close to neutral activation, and we consider it to be in the low arousal category. These findings suggest that predicting dimensions is more challenging than predicting categories.
@@
In summary, the EMOVOME database achieves lower results than other reference databases in the literature. Both RAVDESS and IEMOCAP outperform EMOVOME in emotion categories, and the former also obtains higher valence and arousal prediction results. However, EMOVOME and IEMOCAP exhibit more comparable results in valence and arousal, possibly owing to their text-independent nature (unlike RAVDESS) and the approach used to transform the original labels in IEMOCAP into valence and arousal categories.
VI-C RQ3: What impact do annotator labels exert on the performance of SER models when dealing with challenging natural databases like EMOVOME?
@@
Research on evaluating the influence of annotators’ demographics on SER is limited, but evidence suggests that biases can arise based on their gender, age or educational level [14]. We hypothesize that this bias may be more pronounced when evaluating natural databases (such as EMOVOME) since they do not contain stereotypical emotion (unlike acted databases), and thus can be more challenging to label. In this work, we explore the difference between expert (E) and non-expert (N) annotations, as well as their combination (C) (see Table V and VI). Clinical psychologists, given their professional training and expertise, possess the necessary skills to recognize and interpret emotions, thus can be considered experts in the task.
@@
Surprisingly, the SER models using the expert’s labels achieved the lowest results for both valence and arousal. Non-experts got higher UA values in arousal prediction (58.73%), compared to the combined labels (55.57%) and the expert’s labels (43.57%). For valence, the combined label obtained the highest UA score (61.64%), closely followed by the non-experts (61.36%) and lastly the expert (57.53%). It is worth highlighting that despite the expertise of clinical psychologists in the task, emotions remain highly subjective, and they can be influenced by individual experiences [14]. In fact, an examination of the confusion matrix (see Fig. 3) reveals discernible biases towards specific emotion categories in both valence and arousal results, potentially shaping what the SER models learn. For valence, the model trained on expert labels shows a bias towards positive valence. Conversely, the model trained on non-expert labels tends to misclassify neutral samples as either positive or negative. The combined label model mitigates the bias towards positive expert categories but increases misclassification for negative and neutral categories. In arousal, unlike valence, there’s an unbalanced data distribution (see Table III). Expert labeled 43% of training data as neutral arousal, leading the model to often assign this category to test samples. Non-experts show a bias toward high arousal. Combining labels mitigates expert bias toward the neutral class and reduces non-expert bias toward the positive category. Consequently, the model is trained with fewer low arousal samples, leading to lower accuracy in this category.
@@
Overall, the annotators’ biases may cause differences in UA scores to up to 4% for valence and 15% for arousal. The better performance of the models based on the non-experts and the combined labels may be due to the higher number of annotators included, which may reduce individual biases in their interpretation of emotions.
VI-D RQ4: How does gender impact the fairness of SER model outcomes in EMOVOME compared to reference databases?

### v3

@@ V-C Pre-trained model embeddings and eGeMAPS
VI Results
The cross-validation results in Fig. 4 show the performance across datasets and emotion labels. The UA of the optimal baseline model using eGeMAPS features and machine learning is indicated in gray. Next to this, results for models using embeddings are displayed, along with a third column for the method integrating embeddings and eGeMAPS features (Emb+eGeMAPS). The figure indicates the mean and standard deviation of the UA across the five folds (four for IEMOCAP).
Figure 4: Cross-validation results for the three methods implemented (eGeMAPS, Embeddings and Emb+eGeMAPS) across the different datasets (EMOVOME, IEMOCAP and RAVDESS) and emotion labels (valence, arousal and categories of emotions).
Regarding the baseline eGeMAPS models, the lowest UA values are from the EMOVOME dataset, with 41-44% for valence, 41-51% for arousal and 31.42% for four categories. IEMOCAP achieves 50.20%, 53.21% and 59.50% for valence, arousal and categories, respectively. RAVDESS achieves the highest UA results for valence and arousal, with 57.19% and 62.48%, respectively. For the eight categories, it obtains 45.52%. Overall, SVC was preferred to KNN in most cases.
@@ VI Results
Finally, the Emb+eGeMAPS method obtains similar results to the previous approach. EMOVOME achieves similar values in valence and arousal (53-59% and 52-60% respectively), while the UA in 4-emotion classification decreases (40.22%). A possible reason is that EMOVOME was recorded in-the-wild conditions. Therefore, eGeMAPS features may be affected by microphone quality and background noise [51]. Conversely, IEMOCAP and RAVDESS were recorded in a controlled environment, and the UA values increased around 1-3% compared to the previous approach (except for the emotion categories with IEMOCAP, which slightly decreased). IEMOCAP achieves 62.48%, 60.41% and 68.04% for valence, arousal and categories, respectively. RAVDESS obtains 72.70%, 75.27% and 69.58% for valence, arousal and categories, respectively. Again, unispeech-L is the best option in most cases. However other pre-trained models, with lower results in the previous approach, greatly improved when combined with eGeMAPS features, particularly w2v2-xlsr-53 and w2v2-L-robust.
Using the hyperparameters of the models obtaining the highest UA in cross-validation, new models were trained on the development set and evaluated on the test set, obtaining the results in Table V. The model achieving the highest cross-validation results among the three implemented methods (eGeMAPS, Embeddings and Emb+eGeMAPS) is indicated. The test results follow the trends found in cross-validation, with RAVDESS outperforming IEMOCAP and EMOVOME across the two dimensions (73.54% in valence and 71.94% in arousal). IEMOCAP UA scores are higher than EMOVOME scores in arousal prediction (61.20% vs. 43.57-58.73% respectively), whereas, for valence prediction, IEMOCAP and EMOVOME achieve similar results (60.40% for the former and 57.53-61.64% for the latter). For EMOVOME, we also examined the differences between expert (E) and non-expert (N) annotations, as well as their combination (C). The confusion matrices are represented in Fig. 5.
TABLE V: Test results for 3-class valence and arousal prediction. “PTM” specifies the pre-trained model. For EMOVOME, the rater is indicated: expert (E), non-expert (N), or combined (C).
Dataset Method PTM WA UA
Valence prediction
@@
Arousal prediction
@@
Figure 5: Confusion matrix for the test samples for valence and arousal prediction on EMOVOME, for the different raters.
Considering the emotion categories, test results for the three datasets are presented in Table VI. It includes the model achieving the highest CV results among the three implemented methods. RAVDESS obtains the highest UA score (75.00%), followed by IEMOCAP (69.58%) and EMOVOME (42.58%). To evaluate which emotions are misclassified, Fig. 6 shows the confusion matrix for each dataset.
TABLE VI: Test results for emotion categories prediction.
@@
Figure 7: Evaluation of gender fairness for the three labels and datasets.
VII Discussion
In this study, we addressed the scarcity of spontaneous datasets from real-life environments by developing the EMOVOME dataset—the first public repository featuring genuine emotions from spontaneous conversations in real-life settings. This paper provides a detailed description of EMOVOME and extends access to the scientific community. We evaluated the performance of SER models trained with EMOVOME using various methodologies, and we compared the results with the well-known IEMOCAP dataset, which includes elicited speech, and RAVDESS, comprising acted recordings. Our comparison focused on predicting valence, arousal, and emotion categories. Finally, we also investigated the impact of annotators’ labels and speakers’ genders on model performance. UA was used for comparison since it is more suitable for unbalanced data.
VII-A SER performance
@@
We implemented three methods to create the SER models: the baseline using eGeMAPS features and machine learning, the embeddings from pre-trained models and the integration of embeddings and eGeMAPS features. As shown in Fig. 4, overall, the embedding approach significantly improves model performance compared to the non-transformer baseline, leading to an approximately 10% improvement in unweighted accuracy across all combinations. Unispeech-L consistently showcased better results than other pre-trained models (as shown in [41]), closely followed by hubert-L (which also outperforms wav2vec2 models in previous studies [40]). For the English datasets, in general, the models pre-trained on multiple languages performed worse than those trained on English-only data (as in [40]). Regarding the Emb+eGeMAPS method, it shows similar cross-validation results to the embeddings for the EMOVOME dataset, and UA values slightly increase (max. 3%) for IEMOCAP and RAVDESS in some cases. One potential explanation is that EMOVOME was recorded in natural, uncontrolled conditions, which might impact the reliability of eGeMAPS features, unlike the other two datasets that were conducted in a controlled environment. Unispeech-L is again the top choice in most cases, but notably, other pre-trained models that initially performed lower (e.g. w2v2-xlsr-53 and w2v2-L-robust) show significant improvement when combined with eGeMAPS features.
@@
Considering the test results, the prediction of emotion categories (see Table VI) follow the trend in cross-validation, where EMOVOME obtains the lowest evaluation metric (45.58% UA), followed by IEMOCAP (69.58% UA) and finally RAVDESS (75.00% UA) (even though the first two classify four emotions and the later eight). Although IEMOCAP contains elicited speech, part of the data consists of actors performing scripted dialogues, which could lead us to expect higher classification results. However, IEMOCAP contains utterances whose text content is different throughout the dataset, i.e., it is text-independent (same as EMOVOME). Conversely, RAVDESS is text-dependent, using two fixed sentences. This might have led the models for RAVDESS to prioritize emotional variations over differences in semantic content. Consequently, there is a performance gap in IEMOCAP compared to the other acted datasets, as highlighted in [64]. Nevertheless, our classification results (70.59% WA, 69.58% UA) are comparable to the state of the art for IEMOCAP, which is in the range from 60.0% to 74.3% UA [40], especially considering that we implemented a SI approach, unlike other previous studies. Additionally, we examined the misclassified emotions for each dataset in Fig. 6. For EMOVOME, surprise is not correctly predicted in any case and is mainly confused with happy and neutral emotion, but this category is underrepresented in the data. For IEMOCAP, all emotions are mainly mistaken for the neutral category. For RAVDESS, most emotions are accurately classified ($>$72%).
@@
Regarding the test results for arousal and valence dimensions (see Table V), the SER models for the EMOVOME dataset achieve UA values of 61.64% for valence and 55.57% for arousal, considering the combined label between expert and non-experts. Surprisingly, the IEMOCAP dataset obtains similar values to EMOVOME, 60.04% for valence and 61.20% for arousal. As for RAVDESS, the UA values are 73.54% and 71.94% for valence and arousal, respectively. Initially, one might have anticipated better outcomes in valence, given that earlier studies [65, 40] found that pre-trained models inherently capture linguistic information in the audio, aiding valence prediction. However, in our approach, we utilized pre-trained embeddings solely as feature extractors to obtain speech embeddings without fine-tuning the transformer layers for SER. This step has proven to be fundamental for the models to effectively learn the semantic content [40]. Furthermore, the data used for pre-training significantly impacts the models’ capacity to capture linguistic information, with the inclusion of multi-lingual data adding complexity to the task [40]. These considerations might explain the relatively minor differences observed between valence and arousal predictions. In the case of RAVDESS, the use of fixed semantic content during recording prevented pre-trained models from leveraging text information for valence prediction. Furthermore, an important limitation in both IEMOCAP and RAVDESS is the label transformation applied to obtain the valence and arousal categories, as well as the unbalanced distribution of samples.
@@
In summary, we found that embeddings and Emb+eGeMAPS give similar results for the top-performing pre-trained model, i.e., Unispeech-L. The EMOVOME dataset achieves lower results than other non-natural datasets in the literature. Both RAVDESS and IEMOCAP outperform EMOVOME in emotion categories, and the former also obtains higher valence and arousal prediction results. However, EMOVOME and IEMOCAP exhibit more comparable results in valence and arousal, possibly owing to their text-independent nature (unlike RAVDESS) and the approach used to transform the original labels in IEMOCAP into valence and arousal categories.
VII-B Impact of annotator labels on SER performance
@@
Emotion prediction lacks a definitive ground truth due to its inherently subjective nature. In this study, we addressed this by asking participants to upload audio samples with self-assessed emotional valence (negative, neutral, or positive). To refine the emotional labels, we also incorporated an external evaluation by both experts and non-experts. While research on the impact of annotators’ demographics on SER is limited, existing studies suggests that biases can arise based on their gender, age or educational level [8]. We hypothesized that this bias may be more pronounced when evaluating natural datasets,such as EMOVOME, since they do not contain stereotypical emotion and thus can be more challenging to label. In this work, we explored the difference between expert (E) and non-expert (N) annotations, as well as their combination (C) (see Table V).
@@
Surprisingly, SER models using the expert’s labels achieved the lowest results for both valence and arousal. Non-experts got higher UA values in arousal prediction (58.73%), compared to the combined labels (55.57%) and the expert’s labels (43.57%). For valence, the combined label obtained the highest UA score (61.64%), closely followed by the non-experts (61.36%) and lastly the expert (57.53%). It’s essential to note that while clinical psychologists are experts in the task, emotions remain highly subjective and can be influenced by individual experiences [8]. In fact, the confusion matrix (see Fig. 5) reveals discernible biases towards specific emotion categories in both valence and arousal results, potentially shaping what the SER models learn. For valence, the model trained on expert labels shows a bias towards positive valence. Conversely, the model trained on non-expert labels tends to misclassify neutral samples as either positive or negative. The combined label model mitigates the bias towards positive expert categories but increases misclassification for negative and neutral categories. In arousal, unlike valence, there’s an unbalanced data distribution. Expert labeled 43% of training data as neutral arousal, leading the model to often assign this category to test samples. Non-experts show a bias toward high arousal. Combining labels mitigates expert bias toward the neutral class and reduces non-expert bias toward the positive category. As a result, the model has fewer low arousal samples to train on, lowering accuracy in this category.
@@
Overall, the annotators’ biases may cause differences in UA scores to up to 4% for valence and 15% for arousal. The better performance of the models based on the non-experts and the combined labels may be due to the higher number of annotators included, which may reduce individual biases in their interpretation of emotions. However, this number is still limited compared to the 319 raters in IEMOCAP.
VII-C Fairness evaluation